in april, pat bought a ford mustang with a 5.0-liter engine. the total price, which includes the used car plus sales tax was $17,481.75. if the sales tax was 7.25%, what was the price of the car?
Answer by ikleyn(52781)

.

Let x be the cost of the car.


Then you have this "total money" equation

    x + 0.0725x = 17481.75 dollars,

or

      1.0725x   = 17481.75

      x         = 17481.75/1.0725 = 16300.00  dollars.    ANSWER

Solved.